OpenStudy (anonymous):

2. Solve and graph the inequality |4 – v| < 5. a. Write the inequality as two inequalities without absolute value. b. Solve the inequality and write the solution set. c. Graph the solution on a number line. Please help explain the process in a very simple way. I don't quite get it and I have read over this lesson many times.

OpenStudy (austinl):

Hey there! Welcome to OpenStudy! If you need any help with navigating the site or it's rules just shoot me a message. Now on to your problem, for part (a) do you know what it is asking?

OpenStudy (anonymous):

Not really /:

OpenStudy (jdoe0001):

\(\bf |4 - v| < 5\implies \begin{cases} +(4 - v) < 5\\ \quad \\ \bf -(4 - v) < 5 \end{cases}\)

OpenStudy (austinl):

Okay, lets say you have this inequality, \(|x|<4\) An absolute value is a magnitude, how far is that number from zero. So this means that, that positive number is x from zero and that negative is also x from zero. Make sense? So we can then re-write it as, \(~~~~~~~~x<4\\ -(x)<4\) Now that jdoe has done it, do you see how he did so?

OpenStudy (anonymous):

So basically, part 'a' is asking to write it how it would look if the solution were to be on the positive or negative (two inequalities)? <if that makes sense

OpenStudy (austinl):

Yes, for an absolute value you will always have two equations/inequalities. The positive version of what is in the abs. value as well as a negative version. For part (b) what you need to do is solve for x in each of the inequalities. Do you think you could do that for me?

OpenStudy (anonymous):

I'm a bit stumped at this part. Wouldn't the solution (in my problem) just be " v < 5 " or in your given example problem, wouldn't it just stay x < 4?

OpenStudy (austinl):

Lets look at what jdoe did, which is completely correct. He has, \(\bf |4 - v| < 5\implies \begin{cases} +(4 - v) < 5\\ \quad \\ \bf -(4 - v) < 5 \end{cases}\) So you would need to solve the two inequalities to the right. \(~~~~~~~~4-v<5\\ -(4-v)<5\) Now keep in mind for the second one that you need to distribute that negative.

OpenStudy (anonymous):

I am still really confused. Am I trying to find the value of '4 - v' or just 'v'?

OpenStudy (austinl):

You are trying to solve for v by itself. \(v<\cdot\cdot\cdot\)

OpenStudy (anonymous):

I apologize. I don't know why my brain just won't process this.

OpenStudy (anonymous):

Okay, so it would be 'v < 5' ?

OpenStudy (austinl):

No need, this is clearly new material and new material is always tough. I understand :)

OpenStudy (austinl):

Lets look at the first example shall we? \(4-v<5\) First we would want to get the variable by itself, we would do so by subtracting 4 from both sides. \(-v<5-4\\ -v<1\) Now, we want it to be positive v, so we would divide both sides by negative one. \(\displaystyle \frac{-v}{-1}<\frac{1}{-1}\\ ~\\ v>-1\) And you have to remember to flip the sign whenever you multiply or divide by a negative number. Does this make sense to you?

OpenStudy (anonymous):

Ohhh! That makes sense. So, that limits the solution set down to anything greater than -1 and less than 5 on a number line?

OpenStudy (anonymous):

(Btw sorry it took so long to reply, I was going to draw a number line as an example of how I would graph it but it was difficult)

OpenStudy (austinl):

Um, not quite. Try and solve the second inequality.

OpenStudy (anonymous):

Oh. Now I'm lost again. How exactly do I solve 'v'?

OpenStudy (austinl):

\(-(4-v)<5\) Use the same process that I did before, but be sure to distribute that negative across that sum first :)

OpenStudy (anonymous):

I'm confused on the process that you used before. It says '4 - v < 5' so how did you know to subtract 4 from 5?

OpenStudy (austinl):

Because you need to get terms with a variable on one side and constants on the other. So you can do that by subtracting a term from both sides, or adding. Whatever you do to one side you have to do to the other usually.

OpenStudy (anonymous):

Okay so with the negative inequality would I add 4 to 5?

OpenStudy (austinl):

Yes you would!

OpenStudy (anonymous):

good job helpin @AshB austin :D

OpenStudy (anonymous):

Awesome! So: v < 5+4 v < 9 Then, \[\frac{ x }{ 9 } < \frac{ 1 }{ 9 }\] or would it be \[\frac{ x }{ 9 } > \frac{ 1 }{ 9 }\] ?

OpenStudy (austinl):

You just have, \(v>-1\) \(v<9\) Those are what you have, now do you think you could put those in a solution set?

OpenStudy (anonymous):

Ohh okay. and a solution set is like {0, 1, 2, etc.} correct? << An example, not my answer.

OpenStudy (austinl):

I think it is more like, \((-\infty,-1]\cup[9,\infty)\) I am not quite sure how it would work, I am not so familiar with these.

OpenStudy (anonymous):

So the solution set would contain all numbers and fractions greater than (but not equal to -1) and less than (but not equal to) 9?

OpenStudy (austinl):

Indeed, now do you think you could put this on a number line?

OpenStudy (anonymous):